Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by ritula » Wed Aug 13, 2008 11:06 pm
I know that this one is pretty easy but I am getting answer as 27/28.Pls help


A certain deck of cards contains 2 blue cards, 2 red cards, 2 yellow cards, and 2 green cards. If two cards are randomly drawn from the deck, what is the probability that they will both are not blue?

A. 15/28

B. 1/4

C. 9/16

D. 1/32

E. 1/16
Philosophers have interpreted world in various ways, the point is to change it!
Join the discussion
Source: — Problem Solving |

by amitansu » Wed Aug 13, 2008 11:52 pm
Without blue selection, there are total 6 of them can be chosen.

So for first pick, 6/8
next pick, 5/7

both=> 6/8*5/7=15/28
Join the discussion

by sudhir3127 » Thu Aug 14, 2008 12:54 am
hey i think u probably did

1- (1/28) = 27/28..

even for a second i also did it that way and then saw the answers and realised its not the way..

ur right in ur approach though that...

probably of not blue is 1- ( probability of blue )

but u need to careful when u do this...

heres how u shud do it..

Probability that 1st is blue = 2/8

But 1st will not be blue 1 - 1/4 = 3/4 of the time;
then getting a blue has probability 2/7 (2 blues still left)
since that happens 3/4 of the time,
then: 3/4 * 2/7 = 6/28 =

So add them up: 2/8 + 6/28= 7/28 + 6/28 = 13/28

thus probability of not blue is

1- prob. of blue..

which is 1- 13/28 = 15/28..

though u get an answer this way .. i will still prefer the follow the other way of doing it...

P(1st card not blue) = 6/8 [6 non-blue available out of 8 cards]
P(2nd card not blue given 1st card was not blue) = 5/7 [5 remaining non-blue out of 7]

P(both not blue) = (6/8 )*(5/7) = 15/28

Hope it helps..
Join the discussion

by bluementor » Thu Aug 14, 2008 1:16 am
Hi I think this is a matter of understanding exactly what the question wants, which I'm not sure myself. It could be either one of the following:

1. P(of no blue cards at all) = 1-P(at least 1 blue)

2. P(of no 2 blue cards) = 1 - P(2 blue cards)

From your answers, it seems like its probably (1) that the question is asking. But I'm not sure how do you know that from the question. Could the wording of the question been improved?

BlueMentor
Join the discussion

by pepeprepa » Thu Aug 14, 2008 1:20 am
I read back the question and I do not see 2 ways of understanding it.

"what is the probability that they will both are not blue". For me it is clear, 0 blue in the total.
Join the discussion

Both are not blue vs None is blue

by Ramp » Wed Feb 25, 2009 5:47 pm
Bringing up an old topic .

Both are not blue = P (None is blue )+ P(one is blue and other not)
= 30/56 + 12/56 = 27/28

It can also be solved by
1- P(both are blue) = 1- (2/56) = 27/28

However from the OA (15/28) the question seems to be interpreted as 'None is blue ' case which is just a subset of 'Both are not blue case'.
I think answer should be 27/28

Feedback pls.
Join the discussion